3.1 x 10^5 kg and 6.5 x 10^3 kg the gravitational attraction between them is 65n how far apart are their centers


1
Expert's answer
2021-05-19T11:04:03-0400

"F=G\frac{m_1m_2}{r^2}\to r=\sqrt{\frac{Gm_1m_2}{F}}=\sqrt{\frac{6.67\cdot 10^{-11}\cdot 3.1\cdot10^5\cdot6.5\cdot10^3}{65}}=0.0455\ (m)" . Answer